OpenStudy (sunshinehaggard):

given a right triangle if tanθ=3/4, what is the length of the side adjacent to ∠θ

OpenStudy (kayders1997):

Do you know what tan is equivalent to for the ratio?

OpenStudy (sunshinehaggard):

this is the question it gave me and here are my choices a. 3 b. 4 c. 5 d. 75

OpenStudy (kayders1997):

Okay, do you know what tangent means though?

OpenStudy (kayders1997):

SOH-CAH-TOA

OpenStudy (sunshinehaggard):

yes in a way

OpenStudy (jdoe0001):

recall your SOH CAH TOA \(sin(\theta)=\cfrac{opposite}{hypotenuse} \qquad \qquad % cosine cos(\theta)=\cfrac{adjacent}{hypotenuse} \\ \quad \\ % tangent tan(\theta)=\cfrac{opposite}{{\color{brown}{ adjacent}} }\) it goes well with waffles or pancakes

OpenStudy (kayders1997):

Tangent is opposite/adjacent so what would be your adjeacent side if you have the tan=3/4

OpenStudy (sunshinehaggard):

4?

OpenStudy (kayders1997):

Yes good :)

OpenStudy (sunshinehaggard):

thank you so much

OpenStudy (kayders1997):

You're welcome :) if you need more help you can message me or put my name in the chat for the question you post :)

OpenStudy (kayders1997):

If you would like :)